Tom sells shoes part time at a department store. He can earn either $500 per month plus 4% commission of his total sales, or $400 per month plus 5% commission on his total sales.
What is the total price of the shoes Tom needs to sell to earn the same income from each pay scale? Which is the better offer?

Found 2 solutions by stanbon, mananth:
Answer by stanbon(75887)

Tom sells shoes part time at a department store.
He can earn either $500 per month plus 4% commission of his total sales
S(x) = 500 + 0.04x where x is his sales for the month.
-----------------------------
Or $400 per month plus 5% commission on his total sales.
T(x) = 400 + 0.05x
----------
What is the total price of the shoes Tom needs to sell to earn the same income from each pay scale?
Solve: S(x) = T(x)
500+0.04x = 400 + 0.05x
100 = 0.01x
x = $1000
----------------------
Which is the better offer?
500 + 0.04x R 400+0.5x
100 R 0.01x
--
S is better than T if sales are less than $1000
T is better than S if sales are greater than $1000
=================
Cheers,
Stan H.
========================

Answer by mananth(16949)

Let x be the total sales
Option I 500 + 0.04 x
Option II 400 + 0.05 x
Option I = Option II
500+0.04x=400+0.05 x
500+0.04x=400+0.05x
500-400=0.05x-0.04x
100 = 0.01 x
10000 = x
CHECK
$10000 should be total sales
Option I500 + 0.04 * 10000= 900
Option II400 + 0.05 * 10000= 900

Option II is better offer . the income is more than that of option I when he sells for more than $10,000